# Counting Distinct Elements in the Turnstile Model with Differential Privacy under Continual Observation

### v1

@@ 2 Additional background on differential privacy
In this section we describe basic results on differential privacy used to obtain our theorems.
Lemma 2.1 (Post-processing [28, 11]).
@@
Every $(\varepsilon,\delta)$-DP algorithm $\mathcal{A}$ is $\left(\ell\varepsilon,\delta^{\prime}\right)$-DP for groups of size $\ell$, where $\delta^{\prime}=\delta\frac{e^{\ell\varepsilon}-1}{e^{\varepsilon}-1}$; that is, for all datasets $y,y^{\prime}$ such that $\|y-y^{\prime}\|_{0}\leq\ell$, it holds $\mathcal{A}(y)\approx_{\ell\varepsilon,\delta^{\prime}}\mathcal{A}(y)$.
@@ 2.1 Preliminaries on zero-concentrated differential privacy (zCDP)
This section describes zero-concentrated differential privacy (zCDP), a variant of differential privacy that is less stringent than pure differential privacy, but more stringent than approximate differential privacy. This notion of privacy provides tight bounds for the Gaussian mechanism and cleaner and tighter bounds for composition. In contrast to $(\varepsilon,\delta)$-differential privacy, zCDP requires output distributions on all pairs of neighboring datasets to be $\rho$-close (Definition 2.5) instead of $(\varepsilon,\delta)$-indistinguishable.
@@
The Gaussian mechanism, defined next, is used in Section 3. It privately estimates a real-valued function on a database by adding Gaussian noise to the value of the function.
@@
Let $f:\mathcal{Y}\rightarrow\mathbb{R}^{k}$ be a function. Its $\ell_{2}$-sensitivity is defined as $\max_{\text{neighbors }y,y^{\prime}\in\mathcal{Y}}\|f(y)-f(y^{\prime})\|_{2}.$ To define $\ell_{1}$-sensitivity, we replace the $\ell_{2}$ norm with the $\ell_{1}$ norm.
@@ 3 Item-level private mechanisms for $\mathsf{CountDistinct}$
In this section we prove a version of Theorem 1.4 with zero concentrated differential privacy (zCDP). This notion of privacy provides tight bounds for the Gaussian mechanism and cleaner and tighter bounds for composition. We start by proving Theorem 3.1, and explain in Section 3.3 how this can be used to prove Theorem 1.4.
@@
Fix $\rho\in(0,1]$ and sufficiently large $T\in\mathbb{N}$. Then, there exists a $\rho$-item-level-zCDP mechanism for $\mathsf{CountDistinct}$, that for all turnstile streams $x$ of length $T$ is $\alpha$-accurate where
@@
In Section 3.1, we describe a modification to the binary tree mechanism which, when analysed carefully, provides the desired error guarantees—but only if the maximum flippancy of the stream is known upfront. In Section 3.2, we use this mechanism, in conjunction with a method for adaptively estimating the flippancy bound, to obtain our item-level-DP mechanism for $\mathsf{CountDistinct}$.
@@ 3.1 Enforcing a given flippancy bound $w$
When a flippancy upper bound $w$ is given upfront, we leverage the structure of the binary tree mechanism to privately output the number of distinct elements at each time $t\in[T]$, where $T$ is the stream length. The mechanism and its error guarantees are presented in Algorithm 1 and Theorem 3.2, respectively. To provide intuition, we first describe the mechanism when it is run on streams with maximum flippancy at most $w$. We then discuss a modification that ensures privacy of the mechanism for all streams regardless of maximum flippancy.
Algorithm 1 stores vectors $\tilde{f_{{u}}}\in\{0,1\}^{T}$ for all elements $u\in\mathcal{U}$ that appear in the stream. For streams with maximum flippancy at most $w$, the vector $\tilde{f_{{u}}}$ is equal to the existence vector $f_{u}$. In this case, by Definition 1.1, the number of distinct elements at timestep $t\in[T]$ equals $\sum_{u\in\mathcal{U}}\tilde{f_{{u}}}[t]$. The mechanism therefore outputs values $\sum_{u\in\mathcal{U}}\tilde{f_{{u}}}[t]$ with noise added according to the binary tree mechanism, with privacy parameter $\approx\rho/w$ (see Definition 3.4)—that is, with noise scaled up by a factor of $\sqrt{w}$. The accuracy of this mechanism follows from that of the binary tree mechanism.
However, if the mechanism computed $f_{u}$ instead of $\tilde{f_{{u}}}$, it would not be private for streams with maximum flippancy greater than $w$, since it adds noise that scales according to $w$. To see this, consider item-neighbours $x,x^{\prime}\in\mathcal{U}_{\pm}^{T}$ where $x$ has maximum flippancy $w^{*}>w$: The vectors $\mathsf{CountDistinct}(x)$ and $\mathsf{CountDistinct}(x^{\prime})$ may differ in as many as $\theta(w^{*})$ indices. To provide privacy for such streams, the mechanism simply “truncates” the vector $f_{u}\in\{0,1\}^{T}$ to obtain $\tilde{f_{{u}}}[t]=0$ for all $t\geq t^{*}$ if the flippancy of $u$ in $x[1:t^{*}]$ exceeds $w$. This corresponds to running the naive version of the mechanism (that uses $f_{u}$ instead of $\tilde{f_{{u}}}$) on a “truncated” version of the stream $x$, where all elements in $x$ whose flippancy exceeds $w$ are ignored. (Note that the computation of $\tilde{f_{{u}}}$ can be done online since $\tilde{f_{{u}}}[t]$ depends only on $x[1:t]$.) With a careful analysis of the value stored in each node of the binary-tree, we are able to show that this mechanism is $\rho$-item-level-zCDP for all streams, however, it loses accuracy for streams with many high flippancy elements. In Section 3.2, we leverage this mechanism to provide estimates of $\mathsf{CountDistinct}$ that are both accurate and private for all streams.
@@
Fix $\rho\in(0,1]$, sufficiently large $T\in\mathbb{N}$, and $w\leq T$. There exists a mechanism for $\mathsf{CountDistinct}$ for turnstile streams that is $\rho$-item-level-zCDP for all input streams of length $T$, and $\alpha$-accurate only for streams (of length $T$) with maximum flippancy at most $w$, where $\alpha=O{\left({\frac{\sqrt{w}\log T+\log^{3}T}{\sqrt{\rho}}}\right)}.$
@@
1:Input: Time horizon $T\in\mathbb{N}$, privacy parameter $\rho>0$, flippancy upper bound $w>0$, stream $x\in\mathcal{U}_{\pm}^{T}$.
@@
7:     Suppose $x[t]$ is an insertion or deletion of universe element $u$
@@
12:     Output $s[t]=(\sum_{u\in\mathcal{U}_{x}}\tilde{f_{{u}}}[t])+Z[t]$
@@
For $t\in\mathbb{N}$, the dyadic decomposition of the interval $(0,t]$ is a set of at most $\log t+1$ disjoint intervals whose union is $(0,t]$, obtained as follows. Consider the binary expansion of $t$ (which has at most $\log t+1$ bits), and express $t$ as a sum of distinct powers of $2$ ordered from higher to lower powers. Then, the first interval $(0,r]$ will have size equal to the largest power of $2$ in the sum. The second interval will start at $r+1$ and its size will be equal to the second largest power of $2$ in the sum. Similarly, the remaining intervals are defined until all terms in the summation have been exhausted. For example, for $t=11=8+2+1$, the dyadic decomposition of $(0,11]$ is the intervals $(0,8]$, $(8,10]$ and $(10,11]$.
@@
Let $\rho>0$ and $T\in\mathbb{N}$, such that $T$ is a power of 2. The binary tree with $T$ leaves is a complete binary tree with $T$ leaves labeled as follows. Its $T$ leaves are labeled by the intervals $(t-1,t]$ for all $t\in[T]$ and the internal nodes are labeled by intervals obtained from the union of their children’s intervals. More specifically, the binary tree consists of $\log T+1$ levels. A level $\ell\in[0,\log T]$ partitions the interval $(0,T]$ into a set of $\frac{T}{2^{\ell}}$ disjoint intervals, each of length $2^{\ell}$, of the form $((i-1)\cdot 2^{\ell},i\cdot 2^{\ell}]$. The nodes in level $\ell$ are labelled by the intervals in this partition.
@@
We start by reasoning about the privacy of Algorithm 1. It is helpful to think about Algorithm 1 more explicitly in terms of the binary tree mechanism. We define a mechanism $\mathcal{M}^{\prime}$ that explicitly stores noisy values in the nodes of a binary tree, as in the construction of Definition 3.4, and show that the output of Algorithm 1 can be obtained as a post-processing of the output of $\mathcal{M}^{\prime}$.
Assume w.l.o.g. that $T$ is a power of 2; otherwise, one can consider the value $T^{*}=2^{\lceil\log_{2}T\rceil}$. Fix a stream $x$ as the input to Algorithm 1. For all $t\in[T]$, let $F[t]=\sum_{u\in\mathcal{U}}\tilde{f_{{u}}}[t]$, where the vector $\tilde{f_{{u}}}$ is as obtained by the end of running Algorithm 1 with input $x$. (If $u\notin\mathcal{U}_{x}$, set $\tilde{f_{{u}}}=\mathsf{0}^{T}$. Set $F[0]=0$). Define $\mathcal{M}^{\prime}$ so that on input $x$, for each node $(t_{1},t_{2}]$ of the binary tree with $T$ leaves, it outputs $F[t_{2}]-F[t_{1}]+Z_{(t_{1},t_{2}]}$. We show how to obtain the outputs of Algorithm 1 from the outputs of $\mathcal{M}^{\prime}$. For each time step $t\in[T]$ consider the dyadic decomposition of the interval $(0,t]$ into $k$ intervals $(t_{0},t_{1}],(t_{1},t_{2}],\dots,(t_{k-1},t_{k}]$, corresponding to nodes in the binary tree, where $t_{0}=0$, $t_{k}=t$, and $k\leq\log T+1$. The dyadic decomposition is described in Definition 3.3. We can sum the outputs corresponding to the nodes in the dyadic decomposition of $(0,t]$ to obtain
@@
where we use the fact that $Z$ is a binary tree random variable in the last equality (see Definition 3.4). The right-hand side is exactly the $t$-th output of Algorithm 1.
@@
Note that $G_{\ell}[i]$ + $Z_{(2^{\ell}\cdot(i-1),2^{\ell}\cdot i]}$ equals the output of $\mathcal{M}^{\prime}$ for node $(2^{\ell}\cdot(i-1),2^{\ell}\cdot i]$ in the binary tree. Let $G=(G_{0},G_{1}\dots,G_{\log T})$. Mechanism $\mathcal{M}^{\prime}$ corresponds to applying the Gaussian mechanism (Lemma 2.9) to the output vector $G$, since the variables $Z_{(t_{1},t_{2}]}$ corresponding to the nodes $(t_{1},t_{2}]$ of the binary tree are independent. We now bound the $\ell_{2}$-sensitivity of $G$. Let $x^{\prime}$ be an item-neighboring stream of $x$, and let $u\in\mathcal{U}$ be the universe element on which the two streams differ. Define $\tilde{f_{u}^{\prime}}$, $F^{\prime}$, $G_{\ell}^{\prime}$, and $G^{\prime}$ for the stream $x^{\prime}$ analogously to the definitions of $\tilde{f_{{u}}}$, $F$, $G_{\ell}$, and $G$ for stream $x$.
@@
For all item-neigbhoring streams $x$ and $x^{\prime}$ it holds
@@
We first show that for all levels $\ell\in[0,\log T],$ it holds
@@
Fix some $\ell\in[0,\log T]$ and $i\in[\frac{T}{2^{\ell}}]$. Define $i_{1}=(i-1)\cdot 2^{\ell}$ and $i_{2}=i\cdot 2^{\ell}$. First note that since the streams $x$ and $x^{\prime}$ only differ in the occurrences of element $u$, then the value of $G_{\ell}[i]$ differs by at most $2$ for streams $x$ and $x^{\prime}$:
@@
Observe that $G_{\ell}[i]-G_{\ell}^{\prime}[i]\neq 0$ holds only if at least one of the following hold: $\tilde{f_{{u}}}[i_{1}]\neq\tilde{f_{{u}}}[i_{2}]$ or $\tilde{f_{u}^{\prime}}[i_{1}]\neq\tilde{f_{u}^{\prime}}[i_{2}]$. Define the flippancy of a vector $a\in\mathbb{R}^{T}$, denoted $\mathsf{flip}(a)$, as the number of pairs of adjacent entries of $a$ with different values. The condition $\tilde{f_{{u}}}[i_{1}]\neq\tilde{f_{{u}}}[i_{2}]$ implies that a “flip” occurs in the vector $\tilde{f_{{u}}}$ between indices $i_{1}$ and $i_{2}$. The same holds for $\tilde{f_{u}^{\prime}}$. By the design of Algorithm 1 (and consequently $\mathcal{M}^{\prime}$), $\mathsf{flip}(\tilde{f_{{u}}})\leq w$ and $\mathsf{flip}(\tilde{f_{u}^{\prime}})\leq w$. Additionally, all intervals $(i_{1},i_{2}]$ for a fixed $\ell$ are disjoint. Hence, the number of intervals $i\in[\frac{T}{2^{\ell}}]$ such that $G_{\ell}[i]\neq G^{\prime}_{\ell}[i]$ is at most $2w$. Combining this fact with Equation (3), we obtain the following upper bound on the $\ell_{2}$-sensitivity of $G_{\ell}$ for all levels $\ell\in[0,\log T]$:
@@
With this, we obtain
@@
Recall that mechanism $\mathcal{M}^{\prime}$ corresponds to applying the Gaussian mechanism (Lemma 2.9) to the output vector $G$. By the $\ell_{2}$-sensitivity bound for $G$ (Lemma 3.5), and the privacy of the Gaussian mechanism (Lemma 2.9), we obtain that $\mathcal{M}^{\prime}$ is $28w(\log T+1)\rho^{\prime}/$ 2-zCDP, where $\rho^{\prime}$ is chosen in Step 3 of Algorithm 1. After substituting with the value of $\rho^{\prime}$, it follows that $\mathcal{M}^{\prime}$ (and hence, Algorithm 1) are $\rho$-item-level-zCDP.
Next, we analyze the accuracy of Algorithm 1. Suppose the input stream $x$ has maximum flippancy at most $w$. Then the variables $\tilde{f_{{u}}}$ from Algorithm 1 with input stream $x$ satisfy $\tilde{f_{{u}}}=f_{u}(x)$. Recall that $\mathsf{CountDistinct}(x)\in\mathbb{R}^{T}$ denotes the vector of distinct counts for $x$. Then $\mathsf{CountDistinct}(x)=\sum_{i\in\mathcal{U}}f_{u}(x)=\sum_{i\in\mathcal{U}}\tilde{f_{{u}}}(x)=s-Z$, where $s$ is the vector of outputs of Algorithm 1 defined in Step 12. As a result, $\text{\sf ERR}_{\mathsf{CountDistinct}}(x,s)=\max_{i\in[T]}|Z[t]|$. Each $Z[t]$ is a sum of at most $\log T+1$ independent Gaussian random variables distributed as $\mathcal{N}(0,\frac{1}{\rho^{\prime}})$. Therefore, $Z[t]$ is also Gaussian with mean 0 and variance at most $\frac{\log T+1}{\rho^{\prime}}$. We bound the error of our algorithm by standard concentration inequalities for Gaussian random variables. Set $m=\sqrt{16w(\log T+1)^{2}/\rho}$. By Lemma A.2,

### v3

@@ 2 Additional background on differential privacy
In this section, we describe basic results on differential privacy used to obtain our theorems.
We denote by $\mathcal{N}(\mu,\sigma^{2})$ the Gaussian distribution with mean $\mu$ and standard deviation $\sigma$. The Laplace distribution with mean 00 and standard deviation $\sqrt{2}b$ is denoted by $\mathrm{Lap}(b)$.
@@
Next, we generalize the definition of event-neighboring and item-neighboring streams (Definition 1.3).
Definition 2.2 ($\ell$-Neighboring streams).
Let $x,x^{\prime}\in\mathcal{U}_{\pm}^{T}$ be two streams of length $T$. For any $k\in\mathbb{N}$, streams $x$ and $x^{\prime}$ are $\ell$-event-neighbors (respectively, $\ell$-item-neighbors) if one can be obtained from the other in a sequence of $\ell$ steps, where each step replaces a stream with an event-neighboring (respectively, item-neighboring) stream.
@@
Every $(\varepsilon,\delta)$-event-level (or item-level) DP mechanism $\mathcal{M}$ is $\left(\ell\varepsilon,\delta^{\prime}\right)$-event-level (or item-level) DP for groups of size $\ell$, where $\delta^{\prime}=\delta\frac{e^{\ell\varepsilon}-1}{e^{\varepsilon}-1}$; that is, for all $\ell$-event (or $\ell$-item) neighboring data streams $x,x^{\prime}$, it holds that $\mathcal{M}(x)\approx_{\ell\varepsilon,\delta^{\prime}}\mathcal{M}(x^{\prime})$.
@@ 2.1 Preliminaries on zero-concentrated differential privacy (zCDP)
This section describes zero-concentrated differential privacy (zCDP), a variant of differential privacy that is less stringent than pure differential privacy, but more stringent than approximate differential privacy. Using this notion of privacy, one can show tight bounds for the Gaussian mechanism and cleaner and tighter bounds for composition. In contrast to $(\varepsilon,\delta)$-differential privacy, zCDP requires output distributions on all pairs of neighboring datasets to be $\rho$-close (Definition 2.5) instead of $(\varepsilon,\delta)$-indistinguishable.
@@
Lemma 2.8 (Post-processing [27, 11]).
The Gaussian mechanism, defined next, is used in Section 3. It privately estimates a real-valued function on a dataset by adding Gaussian noise to the value of the function.
@@
Let $f:\mathcal{Y}\rightarrow\mathbb{R}^{k}$ be a function. Its $\ell_{2}$-sensitivity is defined as
$\max_{\text{neighbors }y,y^{\prime}\in\mathcal{Y}}\|f(y)-f(y^{\prime})\|_{2}.$
To define $\ell_{1}$-sensitivity, we replace the $\ell_{2}$ norm with the $\ell_{1}$ norm.
@@ 3 Item-level private mechanisms for $\mathsf{CountDistinct}$
In this section, we prove a version of Theorem 1.5 with zero concentrated differential privacy (zCDP). Using this notion of privacy, one can show tight bounds for the Gaussian mechanism and cleaner and tighter bounds for composition. We start by proving Theorem 3.1, and explain in Section 3.3 how this can be used to prove Theorem 1.5.
@@
For all $\rho\in(0,1]$ and sufficiently large $T\in\mathbb{N}$, there exists a $\rho$-item-level-zCDP mechanism for $\mathsf{CountDistinct}$ that is $\alpha$-accurate for all turnstile streams $x$ of length $T$, where
@@
In Section 3.1, we describe a modification to the binary-tree mechanism which, when analyzed carefully, provides the desired error guarantees—but only if the maximum flippancy of the stream is known upfront. In Section 3.2, we use this mechanism, in conjunction with a method for adaptively estimating the flippancy bound, to obtain our item-level-DP mechanism for $\mathsf{CountDistinct}$.
@@ 3.1 Enforcing a given flippancy bound $w$
When a flippancy upper bound $w$ is given upfront, we leverage the structure of the binary-tree mechanism to privately output the number of distinct elements at each time $t\in[T]$, where $T$ is the stream length. The mechanism and its error guarantees are presented in Algorithm 1 and Theorem 3.2, respectively. To provide intuition, we first describe the mechanism when it is run on streams with maximum flippancy at most $w$. We then discuss a modification that ensures privacy of the mechanism for all streams regardless of maximum flippancy.
Algorithm 1 stores vectors $\tilde{f_{{u}}}\in\{0,1\}^{T}$ for all elements $u\in\mathcal{U}$ that appear in the stream. For streams with maximum flippancy at most $w$, the vector $\tilde{f_{{u}}}$ is equal to the existence vector $f_{u}$. In this case, by Definition 1.1, the number of distinct elements at timestep $t\in[T]$ equals $\sum_{u\in\mathcal{U}}\tilde{f_{{u}}}[t]$. The mechanism outputs values $\sum_{u\in\mathcal{U}}\tilde{f_{{u}}}[t]$ with Gaussian noise added according to the binary-tree mechanism, with privacy parameter $\approx\rho/w$ (see Definition 3.4)—that is, with noise scaled up by a factor of $\approx\sqrt{w/\rho}$. The accuracy of this mechanism follows from that of the binary-tree mechanism.
However, if the mechanism computed $f_{u}$ instead of $\tilde{f_{{u}}}$, it would not be private for streams with maximum flippancy greater than $w$, since it adds noise that scales according to $w$. That is because for every stream $x\in\mathcal{U}_{\pm}^{T}$ with maximum flippancy $w_{x}>w$ there exists a neighboring stream $x^{\prime}$ such that the vectors $\mathsf{CountDistinct}(x)$ and $\mathsf{CountDistinct}(x^{\prime})$ differ in as many as $\Theta(w_{x})$ indices. To provide privacy for such streams, the mechanism simply “truncates” the vector $f_{u}\in\{0,1\}^{T}$ to obtain $\tilde{f_{{u}}}[t]=0$ for all $t\geq t^{*}$ if the flippancy of $u$ in $x[1:t^{*}]$ exceeds $w$. This corresponds to running the naive version of the mechanism (that uses $f_{u}$ instead of $\tilde{f_{{u}}}$) on the “truncated” version of the stream $x$, where elements in $x$ are ignored after their flippancy exceeds $w$. (Note that the computation of $\tilde{f_{{u}}}$ can be done online since $\tilde{f_{{u}}}[t]$ depends only on $x[1:t]$.) With a careful analysis of the value stored in each node of the binary tree, we are able to show that this mechanism is $\rho$-item-level-zCDP for all streams, however, it loses accuracy for streams with many high flippancy elements. In Section 3.2, we leverage this mechanism to provide estimates of $\mathsf{CountDistinct}$ that are both accurate and private for all streams.
@@
Fix $\rho\in(0,1]$, sufficiently large $T\in\mathbb{N}$, and $w\leq T$. Algorithm 1 is a mechanism for $\mathsf{CountDistinct}$ for turnstile streams that is $\rho$-item-level-zCDP for all input streams of length $T$, and $\alpha$-accurate for streams of length $T$ with maximum flippancy at most $w$, where $\alpha=O{\left({\frac{\sqrt{w}\log T+\log^{3}T}{\sqrt{\rho}}}\right)}.$
@@
1:Input: Time horizon $T\in\mathbb{N}$, privacy parameter $\rho>0$, flippancy upper bound $w>0$, stream $x\in\mathcal{U}_{\pm}^{T}$
@@
7:     Suppose $x[t]$ is an insertion or deletion of a universe element $u$
@@
12:     Return $s[t]=(\sum_{u\in\mathcal{U}_{x}}\tilde{f_{{u}}}[t])+Z[t]$
@@
For $t\in\mathbb{N}$, the dyadic decomposition of the interval $(0,t]$ is a set of at most $\log t+1$ disjoint intervals whose union is $(0,t]$, obtained as follows. Consider the binary representation of $t$ (which has at most $\log t+1$ bits), and express $t$ as a sum of distinct powers of $2$. Then, the first interval is $(0,r]$, where $r$ is the largest power of $2$ in the sum. The second interval starts at $r+1$ and its size is the second largest power of $2$ in the sum. The remaining intervals are defined similarly for all remaining summands. For example, for $t=11=8+2+1$, the dyadic decomposition of $(0,11]$ is the intervals $(0,8]$, $(8,10]$ and $(10,11]$.
@@
Let $\rho>0$ be a privacy parameter and $T\in\mathbb{N}$ be a power of 2. Consider a complete binary tree with $T$ leaves whose nodes are labeled as follows. The $T$ leaves are labeled by the intervals $(t-1,t]$ for all $t\in[T]$ and the internal nodes are labeled by intervals obtained from the union of their children’s intervals. Specifically, the binary tree consists of $\log T+1$ levels. A level $\ell\in[0,\log T]$ partitions the interval $(0,T]$ into a set of $\frac{T}{2^{\ell}}$ disjoint intervals, each of length $2^{\ell}$, of the form $((i-1)\cdot 2^{\ell},i\cdot 2^{\ell}]$. The nodes in level $\ell$ are labelled by the intervals in this partition.
@@
We start by reasoning about the privacy of Algorithm 1. It is helpful to think about Algorithm 1 more explicitly in terms of the binary tree mechanism. We define a mechanism $\mathcal{M}^{\prime}$ that returns noisy values for all nodes of the binary tree from Definition 3.4 and show that the output of Algorithm 1 can be obtained by post-processing the output of $\mathcal{M}^{\prime}$.
Assume w.l.o.g. that $T$ is a power of 2; otherwise, consider the value $2^{\lceil\log_{2}T\rceil}$ instead. Fix a stream $x$ as the input to Algorithm 1. For all $t\in[T]$, let $F[t]=\sum_{u\in\mathcal{U}}\tilde{f_{{u}}}[t]$, where the vector $\tilde{f_{{u}}}$ is obtained by the end of running Algorithm 1 with input $x$. (If $u\notin\mathcal{U}_{x}$, set $\tilde{f_{{u}}}=\mathsf{0}^{T}$. Set $F[0]=0$). Define $\mathcal{M}^{\prime}$ so that on input $x$, for each node $(t_{1},t_{2}]$ of the binary tree with $T$ leaves, it outputs $F[t_{2}]-F[t_{1}]+Z_{(t_{1},t_{2}]}$.
We show how to obtain the outputs of Algorithm 1 from the outputs of $\mathcal{M}^{\prime}$. For each time step $t\in[T]$ consider the dyadic decomposition of the interval $(0,t]$ into $k$ intervals $(t_{0},t_{1}],(t_{1},t_{2}],\dots,(t_{k-1},t_{k}]$, corresponding to nodes in the binary tree, where $t_{0}=0$, $t_{k}=t$, and $k\leq\log T+1$. Add the outputs corresponding to the nodes in the dyadic decomposition of $(0,t]$ to obtain
@@
where the last equality holds because $Z$ is a binary-tree random variable (see Definition 3.4). The right-hand side is exactly the $t$-th output of Algorithm 1.
@@
The random variable $G_{\ell}[i]$ + $Z_{(2^{\ell}\cdot(i-1),2^{\ell}\cdot i]}$ equals the output of $\mathcal{M}^{\prime}$ for node $(2^{\ell}\cdot(i-1),2^{\ell}\cdot i]$ in the binary tree. Let $G=(G_{0},G_{1}\dots,G_{\log T})$. Mechanism $\mathcal{M}^{\prime}$ corresponds to applying the Gaussian mechanism (Lemma 2.10) to the output vector $G$, since the variables $Z_{(t_{1},t_{2}]}$ corresponding to the nodes $(t_{1},t_{2}]$ of the binary tree are independent. We now bound the $\ell_{2}$-sensitivity of $G$. Let $x^{\prime}$ be an item-neighboring stream of $x$, and let $u\in\mathcal{U}$ be the universe element on which the two streams differ. Define $\tilde{f_{u}^{\prime}}$, $F^{\prime}$, $G_{\ell}^{\prime}$, and $G^{\prime}$ for the stream $x^{\prime}$ analogously to the definitions of $\tilde{f_{{u}}}$, $F$, $G_{\ell}$, and $G$ for stream $x$.
@@
For all item-neighboring streams $x$ and $x^{\prime}$,
@@
We first show that for all levels $\ell\in[0,\log T]$,
@@
Fix some $\ell\in[0,\log T]$ and $i\in[\frac{T}{2^{\ell}}]$. Define $i_{1}=(i-1)\cdot 2^{\ell}$ and $i_{2}=i\cdot 2^{\ell}$. Since the streams $x$ and $x^{\prime}$ only differ in the occurrences of element $u$, the values $G_{\ell}[i]$ and $G_{\ell}^{\prime}[i]$ differ by at most $2$:
@@
Observe that $G_{\ell}[i]-G_{\ell}^{\prime}[i]\neq 0$ only if at least one of the following hold: $\tilde{f_{{u}}}[i_{1}]\neq\tilde{f_{{u}}}[i_{2}]$ or $\tilde{f_{u}^{\prime}}[i_{1}]\neq\tilde{f_{u}^{\prime}}[i_{2}]$. Define the flippancy of a vector $a\in\mathbb{R}^{T}$, denoted $\mathsf{flip}(a)$, as the number of pairs of adjacent entries of $a$ with different values. The condition $\tilde{f_{{u}}}[i_{1}]\neq\tilde{f_{{u}}}[i_{2}]$ implies that a “flip” occurs in the vector $\tilde{f_{{u}}}$ between indices $i_{1}$ and $i_{2}$. The same holds for $\tilde{f_{u}^{\prime}}$. By the design of Algorithm 1 (and consequently $\mathcal{M}^{\prime}$), $\mathsf{flip}(\tilde{f_{{u}}})\leq w$ and $\mathsf{flip}(\tilde{f_{u}^{\prime}})\leq w$. Additionally, all intervals $(i_{1},i_{2}]$ for a fixed $\ell$ are disjoint. Hence, the number of intervals $i\in[\frac{T}{2^{\ell}}]$ such that $G_{\ell}[i]\neq G^{\prime}_{\ell}[i]$ is at most $2w$. Combining this fact with Equation (3), we obtain the following upper bound on the $\ell_{2}$-sensitivity of $G_{\ell}$ for all levels $\ell\in[0,\log T]$:
@@
Combining the equalities for all levels, we obtain
@@
Recall that mechanism $\mathcal{M}^{\prime}$ corresponds to applying the Gaussian mechanism to the output vector $G$. By the $\ell_{2}$-sensitivity bound for $G$ (Lemma 3.5), and the privacy of the Gaussian mechanism (Lemma 2.10), we obtain that $\mathcal{M}^{\prime}$ is $(8w(\log T+1)\rho^{\prime}/2)$-zCDP, where $\rho^{\prime}$ is chosen in Step 3 of Algorithm 1. By the choice of $\rho^{\prime}$, mechanism $\mathcal{M}^{\prime}$ (and hence, Algorithm 1) is $\rho$-item-level-zCDP.
Next, we analyze the accuracy of Algorithm 1. Suppose the input stream $x$ has maximum flippancy at most $w$. Then the variables $\tilde{f_{{u}}}$ from Algorithm 1 with input stream $x$ satisfy $\tilde{f_{{u}}}=f_{u}(x)$. Recall that $\mathsf{CountDistinct}(x)\in\mathbb{R}^{T}$ denotes the vector of distinct counts for $x$. Then $\mathsf{CountDistinct}(x)=\sum_{i\in\mathcal{U}}f_{u}(x)=\sum_{i\in\mathcal{U}}\tilde{f_{{u}}}(x)=s-Z$, where $s$ is the vector of outputs of Algorithm 1 defined in Step 12. As a result, $\text{\sf ERR}_{\mathsf{CountDistinct}}(x,s)=\max_{i\in[T]}|Z[t]|$. Each $Z[t]$ is the sum of at most $\log T+1$ independent Gaussian random variables distributed as $\mathcal{N}(0,\frac{1}{\rho^{\prime}})$. Therefore, $Z[t]$ is also Gaussian with mean 0 and variance at most $\frac{\log T+1}{\rho^{\prime}}$. We bound the error of our algorithm by standard concentration inequalities for Gaussian random variables. Set $m=\sqrt{16w(\log T+1)^{2}/\rho}$. By Lemma C.2,